# FlowKV: Enhancing Multi-Turn Conversational Coherence in LLMs via Isolated Key-Value Cache Management

### v1

@@ Abstract
FlowKV: Enhancing Multi-Turn Conversational Coherence in LLMs via Isolated Key-Value Cache Management
Xiang LIU        Hong CHEN        Xuming HU        Xiaowen CHU The Hong Kong University of Science and Technology(Guangzhou) {xliu886,hchen763}@connect.hkust-gz.edu.cn {xuminghu, xwchu}@hkust-gz.edu.cn
**footnotetext: Equal Contribution.$\dagger$\$$\dagger$\$footnotetext: Corresponding Author.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success in a wide range of natural language understanding and generation tasks, powering applications from open-domain dialogue systems to complex instruction following (Raffel et al., 2020; Brown et al., 2020; Chowdhery et al., 2022; Tay et al., 2022; Touvron et al., 2023a, b; Dubey et al., 2024; Guo et al., 2025). However, as these models are increasingly deployed in multi-turn conversational scenarios, new challenges emerge regarding their efficiency and ability to maintain coherent, contextually relevant responses over extended interactions.
Figure 1: Performance comparison of different KV Cache compression methods on multi-turn benchmark PrefEval. # SKV: SnapKV, # EA: ExpectedAttention.
@@
This challenge of efficiently managing extensive conversational history, so critical for maintaining coherence as highlighted by the attention patterns, is primarily rooted in the operational characteristics of the KV Cache within the attention mechanism. While the KV Cache enables efficient autoregressive generation by storing intermediate representations, its memory footprint grows linearly with the length of the conversation history (Xiao et al., 2023). This not only imposes significant computational and storage costs but also limits the practical deployment of LLMs in real-world, resource-constrained environments.
To address these challenges, a variety of KV Cache eviction strategies have been proposed (Xiao et al., 2023; Zhang et al., 2023; Liu et al., 2025b; Li et al., 2024b; Liu et al., 2025a; Cai et al., 2024; Yang et al., 2024b). Despite their promise, the effectiveness of these methods in multi-turn dialogue settings remains underexplored. In particular, it is unclear how different compression strategies impact the model’s ability to follow instructions, retain user preferences, and maintain long-term dependencies across multiple conversational turns.
In this work, we conduct a comprehensive empirical study of KV Cache eviction strategies in multi-turn dialogue scenarios. We benchmark several state-of-the-art techniques—including SnapKV Li et al. (2024b), StreamingLLM Xiao et al. (2023), ExpectedAttention Jegou et al. (2024), and ChunkKV Liu et al. (2025b)—on two representative datasets: Multi-IF He et al. (2024), which evaluates instruction following across turns, and PrefEval Zhao et al. (2025), which assesses user preference retention. Figure 1 shows these methods’ performance on PrefEval, highlighting the KV cache compression method still have a large gap with the full KV cache. Furthermore, we introduce FlowKV, a multi-turn isolation mechanism designed to mitigate information loss and context forgetting under aggressive compression. This training-free approach is compatible with any existing KV Cache compression method and demonstrates significant improvements, for instance, boosting the average Instruction Following Rate (IFR) by over 20% on Multi-IF and increasing user preference adherence on PrefEval from as low as 10.90% to 75.40% with LLaMA-3.1-8B.
@@ 2.1 Preliminaries: KV Cache Dynamics in Multi-Turn Conversations
The KV cache is a fundamental component in Transformer-based LLMs (Vaswani et al., 2017), designed to enhance the efficiency of autoregressive token generation. It stores the computed key and value vectors for all preceding tokens in a sequence, thereby obviating the need for their recomputation at each subsequent generation step.
@@
For any subsequent Turn $t$ ($t>1$), let $H_{t-1}$ be the complete conversational history up to the end of turn $t-1$:
@@
In contrast to the Full KV Cache, many traditional KV cache eviction strategies aim to reduce memory by compressing the conversational history accumulated up to the current turn, before processing the current user query. A common approach, similar to mechanisms found in frameworks like NVIDIA’s kvpress (Jegou et al., 2024), while leaving the KV cache of the current turn’s query uncompressed for better performance. Let $\mathcal{F}(\cdot)$ denote a lossy KV cache compression.
@@
This pattern continues. After $T$ turns, the KV cache corresponding to the initial system prompt, $P_{sys}$, would have been subjected to $T$ nested compression operations on the path it contributes to the final compressed history. And $KV(Q_{i})$ and $KV(R_{i})$ for $i\leq t$ will be compressed $T-i$ times.
@@ 3.1 Experimental Setups
To evaluate the performance of LLMs under multi-turn conversation with different KV Cache compression strategies, we assess the Multi-IF He et al. (2024) and the PrefEval Zhao et al. (2025) datasets. Multi-IF focuses on evaluating the continuous instruction-following ability of LLMs across multiple turns of conversation, and directly identifies the influence of compression strategies on long-range context dependence and instruction execution accuracy. Besides, PrefEval provides diverse and challenging conversation prompts with hidden user preferences. It can be used to evaluate the impact of compression on the model’s ability to maintain conversation coherence and follow user preferences in complex interactions.
@@
We conduct experiments on two LLMs, including LLaMA-3.1-8B-Instruct Grattafiori et al. (2024) and Qwen-2.5-7B-Instruct Yang et al. (2024a).
@@
We selected the following KV cache compression methods for a comprehensive investigation of their effects: SnapKV Li et al. (2024b), StreamingLLM Xiao et al. (2023), ExpectedAttention Jegou et al. (2024), ChunkKV Liu et al. (2025b) and FlowKV. For below experiments, baseline means the method without FlowKV.
@@
We apply the kvpress Jegou et al. (2024) library to load models and KV Cache compression methods and evaluate their performance using a server with NVIDIA A40 GPUs. All experiments are conducted three times and the average results are reported.
3.2 Performance Comparison on Multi-IF
@@ 3.3 Performance Comparison on PrefEval
To further investigate the performance of FlowKV under varying compression strengths and its compatibility with different KV cache management methods, we plotted the user preference following rate as a function of the compression ratio. As shown in Figure 6, we present results for LLaMA-3.1-8B-Instruct with SnapKV and ExpectedAttention, both with and without FlowKV, on the PrefEval benchmark. First, by comparing the full KV cache and the baselines without FlowKV (lines without markers), we observe a substantial performance drop even at low compression ratios. This finding highlights that user preference following in multi-turn conversations is highly sensitive to the compression ratio. When FlowKV is incorporated, the performance improves markedly. At lower compression ratios (0.1–0.4), FlowKV is able to recover performance to a level close to that of the full KV cache. At higher compression ratios (0.5–0.9), FlowKV continues to provide significant improvements, although the gains are somewhat reduced compared to the low compression ratio. These results demonstrate that FlowKV can robustly preserve information flow across multiple dialogue turns, regardless of the compression ratio. Furthermore, our findings reveal that user preference following in multi-turn dialogue scenarios remains a highly challenging task.
3.4 Case Study
To better illustrate the impact of model generation quality across different multi-turn strategies and KV Cache compression methods, we conduct a case study to demonstrate the performance differences of the LLaMA-3.1-8B-Instruct model when processing multi-turn cumulative instructions on the Multi-IF dataset. Figure 5 shows the Full KV Cache setting successfully follows all instructions of each turn, accurately generating content that satisfies format control, style restrictions, stipulated texts, and other complex tasks. However, the Baseline setting leads to a significant performance crash: not only does content duplication occur in the second turn, violating the newly added "p.p.s" instructions, but in the third turn, the "p.p.s" and double quotation instructions fail to be fully executed, exposing severe context forgetting and the loss of processing capabilities for the conversation flow. When using FlowKV multi-turn conversation, the model shows relative adaptability and performance trade-offs. Although the requirement of all-lowercase format accumulated in the first turn is not fully maintained from the second turn, the content generation still maintains relevance and creativity to the topic and successfully executes the "p.p.s" marks and double quotation marks wrapping instructions newly added in the second and third turns, respectively. This indicates that FlowKV has effectively preserved historical information flow also keep the low KV Cache compression ratio. More case study can be found in Appendix C.
@@ 4 Related Works
4.1 Key–Value Cache Optimization Strategies on LLMs
The attention mechanism needs to compute the relationship between the current token and previous tokens when autoregressive generation. LLMs introduce KV Cache to avoid re-computing the Key and Value vectors of previous tokens. However, the size of KV Cache has a linear relationship with the sequence length Bai et al. (2023), and it may increase sharply when handling long context or multi-turn conversation. To alleviate the challenges brought by KV Cache, researchers proposed multiple compression and optimization strategies, which can be roughly divided into quantization, sparsification (pruning and eviction), low-rank approximation, and cross-layer sharing. The token sparsification here identifies and selectively retains the KV Cache corresponding to the most important tokens, discarding the relatively unimportant parts. For instance, H2O Zhang et al. (2023) considers the token with a higher cumulative attention score to be more important, while FastGen Ge et al. (2023) customizes different strategies for heads by analyzing the behavioral patterns of attention heads. StreamingLLM Xiao et al. (2023) retains only some of the initial tokens and a fixed-size window of the most recent tokens based on its discovery that these tokens are crucial for maintaining the model’s performance. In addition, SnapKV Li et al. (2024b) compresses KV caches by selecting key attention head features from a prompt, enabling efficient long-sequence processing with minimal performance loss. ChunkKV Liu et al. (2025b) compresses KV caches by grouping tokens into semantic chunks, preserving contextual meaning while reducing memory usage. Besides, low-rank approximation Saxena et al. (2024) and cross-layer sharing Sun et al. (2024); Brandon et al. (2024); Liu et al. (2024) and others are also important optimization directions.
@@ 4.2 Multi-turn Conversation in LLMs
Multi-turn conversation represents a crucial application scenario for LLMs, simulating human conversation to understand context and generate coherent, relevant responses over successive interaction turns Zhang et al. (2019); Shuster et al. (2022); Thoppilan et al. (2022); Wu et al. (2020). However, as the number of turns increases, LLMs encounter challenges. Current attention is largely focused on Long Context Handling Yi et al. (2024), which involves treating preceding turns as contextual information and managing the growing conversation history, especially given the constraints of models’ limited context windows. Furthermore, Coherence and Consistency Li et al. (2025) pose another challenge, concerning the need to ensure that responses generated in later turns remain consistent with earlier parts of the conversation and avoid contradictions.
@@ Acknowledgments
This work was supported by Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
Appendix A Additional Results
A.1 Efficiency Results
@@
Appendix B Additional Related Work
B.1 Multi-Turn Conversation
Long-term Dependency Sun et al. (2025) is also a significant focus, requiring models to accurately grasp entities, relationships, and user intentions established early in the conversation and generate responses based on this understanding. To address these challenges, integrating Retrieval-Augmented Generation (RAG) Lewis et al. (2020) with LLMs can improve the accuracy and relevance of responses. Additionally, Yi et al. (2024) and Madotto et al. (2019) mention that explicitly modeling the structural information of the conversation can effectively convey inter-turn dependencies and speaker information to the model for reference. Benchmark datasets for multi-turn conversation often evaluate the ability of LLMs to follow instructions or retain information over multiple accumulated turns. For instance, He et al. (2024) introduces the Multi-IF dataset to assess the degree to which different LLMs continuously adhere to user instructions across turns in multi-turn scenarios. Zhao et al. (2025) constructs the PrefEval dataset, containing user preference and query pairs presented in a multi-turn format, designed to evaluate the capability of large LLMs to infer, remember, and follow user preferences within long-context, multi-turn conversations. SCBench (Li et al., 2024a) is a comprehensice benchmark for evaluating long-context, and multi-turn conversation is main subtask.
Appendix C Case Study
@@
Appendix D Evaluation Benchmark
D.1 Dataset Details

### v2

# Multi-Turn Interactions in Large Language Models

@@ Abstract
[…]
Figure 1: KV cache compression on PrefEval. (#SKV: SnapKV, #EA: ExpectedAttention).
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success in a wide range of natural language understanding and generation tasks, powering applications from open-domain dialogue systems to complex instruction following [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, as these models are increasingly deployed in multi-turn conversational scenarios, new challenges emerge regarding their efficiency and ability to maintain coherent, contextually relevant responses over extended interactions.
@@
This challenge of efficiently managing extensive conversational history, so critical for maintaining coherence as highlighted by the attention patterns, is primarily rooted in the operational characteristics of the KV Cache within the attention mechanism. While the KV Cache enables efficient autoregressive generation by storing intermediate representations, its memory footprint grows linearly with the length of the conversation history [11]. This not only imposes significant computational and storage costs but also limits the practical deployment of LLMs in real-world, resource-constrained environments.
To address these challenges, a variety of KV Cache eviction strategies have been proposed [11, 12, 13, 14, 15, 16, 17, 18, 19]. Despite their promise, the effectiveness of these methods in multi-turn dialogue settings remains underexplored. In particular, it is unclear how different compression strategies impact the model’s ability to follow instructions, retain user preferences, and maintain long-term dependencies across multiple conversational turns.
In this work, we conduct a comprehensive empirical study of KV Cache eviction strategies in multi-turn dialogue scenarios. We benchmark several state-of-the-art techniques—including SnapKV [14], StreamingLLM [11], ExpectedAttention [20], and ChunkKV [13]—on two representative datasets: Multi-IF [21], which evaluates instruction following across turns, and PrefEval [22], which assesses user preference retention. Figure 1 shows these methods’ performance on PrefEval, highlighting the KV cache compression method still have a large gap with the full KV cache. Furthermore, we introduce FlowKV, a multi-turn isolation mechanism designed to mitigate information loss and context forgetting under aggressive compression. This training-free approach is compatible with any existing KV Cache compression method and demonstrates significant improvements, for instance, boosting the average Instruction Following Rate (IFR) by over 20% on Multi-IF and increasing user preference adherence on PrefEval from as low as 10.90% to 75.40% with LLaMA-3.1-8B.
@@ 2.1 Preliminaries: KV Cache Dynamics in Multi-Turn Conversations
The KV cache is a fundamental component in Transformer-based LLMs [23], designed to enhance the efficiency of autoregressive token generation. It stores the computed key and value vectors for all preceding tokens in a sequence, thereby obviating the need for their recomputation at each subsequent generation step.
@@
For any subsequent Turn $t$ ($t>1$), let $H_{t-1}$ be the complete conversational history up to the end of turn t1t-1:
@@
In contrast to the Full KV Cache, many traditional KV cache eviction strategies aim to reduce memory by compressing the conversational history accumulated up to the current turn, before processing the current user query. A common approach, similar to mechanisms found in frameworks like NVIDIA’s kvpress [20], while leaving the KV cache of the current turn’s query uncompressed for better performance.
@@
This pattern continues. After $T$ turns, the KV cache corresponding to the initial system prompt, $P_{sys}$, would have been subjected to $T$ nested compression operations on the path it contributes to the final compressed history. And $KV(Q_{i})$ and $KV(R_{i})$ for $i\leq t$ will be compressed TiT-i times.
@@ 3.1 Experimental Setups
To evaluate the performance of LLMs under multi-turn conversation with different KV Cache compression strategies, we assess the Multi-IF [21] and the PrefEval [22] datasets. Multi-IF focuses on evaluating the continuous instruction-following ability of LLMs across multiple turns of conversation, and directly identifies the influence of compression strategies on long-range context dependence and instruction execution accuracy. Besides, PrefEval provides diverse and challenging conversation prompts with hidden user preferences. It can be used to evaluate the impact of compression on the model’s ability to maintain conversation coherence and follow user preferences in complex interactions.
@@
We conduct experiments on two LLMs, including LLaMA-3.1-8B-Instruct [24] and Qwen-2.5-7B-Instruct [25].
@@
We selected the following KV cache compression methods for a comprehensive investigation of their effects: SnapKV [14], StreamingLLM [11], ExpectedAttention [20], ChunkKV [13] and FlowKV. For below experiments, baseline means the method without FlowKV.
@@
We apply the kvpress [20] library to load models and KV Cache compression methods and evaluate their performance using a server with NVIDIA A40 GPUs. All experiments are conducted three times and the average results are reported.
@@ 3.3 Performance Comparison on PrefEval
To further investigate the performance of FlowKV under varying compression strengths and its compatibility with different KV cache management methods, we plotted the user preference following rate as a function of the compression ratio. As shown in Figure 5, we present results for LLaMA-3.1-8B-Instruct with SnapKV and ExpectedAttention, both with and without FlowKV, on the PrefEval benchmark. First, by comparing the full KV cache and the baselines without FlowKV (lines without markers), we observe a substantial performance drop even at low compression ratios. This finding highlights that user preference following in multi-turn conversations is highly sensitive to the compression ratio.
When FlowKV is incorporated, the performance improves markedly. At lower compression ratios (0.1–0.4), FlowKV is able to recover performance to a level close to that of the full KV cache. At higher compression ratios (0.5–0.9), FlowKV continues to provide significant improvements, although the gains are somewhat reduced compared to the low compression ratio.These results demonstrate that FlowKV can robustly preserve information flow across multiple dialogue turns, regardless of the compression ratio. Furthermore, our findings reveal that user preference following in multi-turn dialogue scenarios remains a highly challenging task.
3.4 Case Study
@@
To better illustrate the impact of model generation quality across different multi-turn strategies and KV Cache compression methods, we conduct a case study to demonstrate the performance differences of the LLaMA-3.1-8B-Instruct model when processing multi-turn cumulative instructions on the Multi-IF dataset. Figure 6 shows the Full KV Cache setting successfully follows all instructions of each turn, accurately generating content that satisfies format control, style restrictions, stipulated texts, and other complex tasks. However, the Baseline setting leads to a significant performance crash: not only does content duplication occur in the second turn, violating the newly added "p.p.s" instructions, but in the third turn, the "p.p.s" and double quotation instructions fail to be fully executed, exposing severe context forgetting and the loss of processing capabilities for the conversation flow. When using FlowKV multi-turn conversation, the model shows relative adaptability and performance trade-offs. Although the requirement of all-lowercase format accumulated in the first turn is not fully maintained from the second turn, the content generation still maintains relevance and creativity to the topic and successfully executes the "p.p.s" marks and double quotation marks wrapping instructions newly added in the second and third turns, respectively. This indicates that FlowKV has effectively preserved historical information flow also keep the low KV Cache compression ratio. More case study can be found in Appendix B.
@@ 4.1 Multi-turn Conversation in LLMs
Multi-turn conversation represents a crucial application scenario for LLMs, simulating human conversation to understand context and generate coherent, relevant responses over successive interaction turns [26, 27, 28, 29]. However, as the number of turns increases, LLMs encounter challenges. Current attention is largely focused on Long Context Handling [30], which involves treating preceding turns as contextual information and managing the growing conversation history, especially given the constraints of models’ limited context windows. Furthermore, Coherence and Consistency [31] pose another challenge, concerning the need to ensure that responses generated in later turns remain consistent with earlier parts of the conversation and avoid contradictions. SCBench [32] is a comprehensive benchmark for evaluating long-context, and multi-turn conversation is main subtask.
@@ Acknowledgments
This work was supported by Guangzhou Municipal Joint Funding Project with Universities and Enterprises under Grant No. 2024A03J0616 and Guangzhou Municipality Big Data Intelligence Key Lab (2023A03J0012); Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
Appendix A Additional Experimental Results
For this section, we conduct additional experiments to further validate the generalization, robustness, and compatibility of FlowKV.
A.1 Long-Context and Reasoning Generalization
To assess FlowKV’s performance on tasks requiring reasoning over long contexts, we conducted new experiments on the Code RepoQA subset of SCBench [32]. This benchmark is particularly challenging, featuring multi-turn interactions with an average input length of 64K tokens where the model must reason about a code repository. The results, presented in Table 3, show that while the task is inherently difficult (as indicated by the low FullKV scores), our FlowKV mechanism consistently and significantly improves the performance of the base compression method (SnapKV), bringing it much closer to the FullKV baseline. This demonstrates that the benefits of preventing repeated information loss are even more pronounced as dialogue history and context length grow.
Table 3: LLaMA-3.1-8B-Instruct Performance on SCBench (Code RepoQA Subset) with SnapKV
Method Compression Ratio Turn 1 Turn 2 Turn 3 Turn 4 Turn 5
FullKV - 2.27% 3.41% 6.82% 1.14% 2.27%
SnapKV 0.1 2.27% 0.00% 0.00% 0.00% 0.00%
+FlowKV 2.27% 3.41% 5.68% 1.14% 1.14%
SnapKV 0.3 1.14% 0.00% 0.00% 0.00% 0.00%
+FlowKV 1.14% 2.27% 3.41% 0.00% 0.00%
SnapKV 0.5 1.14% 0.00% 0.00% 0.00% 0.00%
+FlowKV 1.14% 1.14% 3.41% 0.00% 0.00%
A.2 Performance with Smaller Models
A key point raised was whether FlowKV’s performance gains are tied to the strength of the base model. To investigate this, we conducted new experiments with a smaller model, Qwen2.5-1.5B-Instruct. The results, presented in Table 4 for Multi-IF and Table 5 for PrefEval, show that while the absolute performance is lower than with an 8B model (as expected), FlowKV still provides a substantial relative improvement over the baseline compressor. This aligns with FlowKV’s design motivation: to mitigate the performance degradation caused by repeated compression in multi-turn dialogues, regardless of the base model’s strength.
Table 4: Performance on Multi-IF with Qwen2.5-1.5B-Instruct and SnapKV
Method Compression Ratio IFR R1 IFR R2 IFR R3
FullKV - 42.02% 30.18% 26.23%
SnapKV 0.1 44.13% 14.21% 12.45%
+FlowKV 44.13% 29.80% 26.23%
SnapKV 0.3 43.82% 14.35% 12.88%
+FlowKV 43.82% 30.19% 25.79%
SnapKV 0.5 44.71% 14.46% 12.72%
+FlowKV 44.71% 28.55% 25.94%
Table 5: Performance on PrefEval with Qwen2.5-1.5B-Instruct and SnapKV
Method Compression Ratio ACC
FullKV - 20.00%
SnapKV 0.3 2.20%
+FlowKV 18.70%
SnapKV 0.5 1.90%
+FlowKV 15.20%
A.3 Compatibility with Trainable Compression Methods
FlowKV is a universal, training-free mechanism designed to be orthogonal to the underlying compression algorithm. To demonstrate its compatibility with learnable methods, we conducted an experiment with Activation Beacon [33], a trainable KV Cache compression method. We loaded the pre-trained Beacon-Qwen-2-7B-Instruct model and evaluated it on a subset of 200 samples from the Multi-IF benchmark.
The results in Table 6 show that even with a strong, learnable baseline like Activation Beacon, FlowKV provides significant improvements, particularly in the second turn (R2). This confirms that FlowKV can be effectively combined with trainable compression methods to further mitigate information loss in multi-turn scenarios.
Table 6: Performance on Multi-IF with trainable method (Activation Beacon on Qwen-2-7B-Instruct)
Method Compression Ratio IFR R1 IFR R2 IFR R3
FullKV - 49.01% 26.40% 10.43%
Baseline 0.1 41.38% 25.46% 16.01%
+FlowKV 41.38% 32.15% 12.95%
Baseline 0.5 42.66% 25.49% 16.76%
+FlowKV 42.66% 34.11% 18.49%
A.4 Efficiency Results
@@ A.5 Detailed Multi-Turn Benchmark Results
Appendix B Case Study
@@
Appendix C Theoretical Analysis of Information Loss
We provide a formal analysis to demonstrate why FlowKV’s isolation mechanism is effective at mitigating the information loss that plagues traditional nested compression strategies. We model the process using a simplified information loss model.
C.1 A Simplified Information Loss Model
We can model any lossy KV cache compression function $\mathcal{F}$ as a linear transformation that preserves a fraction of the original information while introducing an error term.
Let a KV cache $C$ be represented as a vector. The action of the compression function $\mathcal{F}$ on $C$ can be simplified as:
$\mathcal{F}(C)=\alpha C+\epsilon$ (14)
Where:
  • $C$ is the original input KV cache vector.
  • $\alpha$ is the information retention factor, where $0\leq\alpha<1$. A value of $\alpha$ closer to 1 indicates more information is preserved.
  • $\epsilon$ is an error/noise vector representing the distortion or information loss introduced during compression.
Our goal is to analyze the degradation of the initial context $C_{0}$ (e.g., the KV cache of the system prompt) after multiple conversational turns.
C.2 Analysis of the Traditional Nested Strategy
In a traditional strategy, the entire history, including previously compressed results, is re-compressed at each turn. Let’s track the state of the original information from $C_{0}$:
After Turn 1, the first compression of $C_{0}$ yields $\mathcal{F}(C_{0})=\alpha C_{0}+\epsilon_{0}$.
After Turn 2, the history containing $\mathcal{F}(C_{0})$ is compressed again. Focusing on the transformation of the $C_{0}$ component:
$\mathcal{F}(\mathcal{F}(C_{0}))=\alpha(\alpha C_{0}+\epsilon_{0})+\epsilon_{1}=\alpha^{2}C_{0}+\alpha\epsilon_{0}+\epsilon_{1}$ (15)
After Turn T, by induction, the state of the information from $C_{0}$, denoted as $C_{0}^{(T)}$, becomes:
$C_{0}^{(T)}=\alpha^{T}C_{0}+\sum_{i=0}^{T-1}\alpha^{i}\epsilon_{T-1-i}$ (16)
Analysis:
  • Signal Decay: The original signal $C_{0}$ is attenuated by a factor of $\alpha^{T}$. Since $\alpha<1$, this term decays exponentially with the number of turns $T$. This formally models the concept of "catastrophic forgetting" of long-term dependencies.
  • Error Accumulation: The total error is a weighted sum of all past errors. Early errors are propagated and compounded through subsequent compression steps.
C.3 Analysis of the FlowKV Isolation Strategy
FlowKV’s core principle is to isolate already-compressed caches from re-compression. Let’s track $C_{0}$ again:
After Turn 1, the initial compression is identical to the traditional strategy: $\mathcal{F}(C_{0})=\alpha C_{0}+\epsilon_{0}$.
For turns $t>1$, according to FlowKV’s mechanism, the compressed block $\mathcal{F}(C_{0})$ is preserved and is not subjected to further compression. Therefore, at any subsequent turn $t\geq 1$, the representation of the original information from $C_{0}$ remains unchanged:
$C_{0}^{(t)}=\alpha C_{0}+\epsilon_{0}$ (17)
Analysis:
  • Signal Decay: The original signal $C_{0}$ is attenuated by a constant factor $\alpha$, regardless of the number of turns $T$.
  • Error Accumulation: The error associated with $C_{0}$ is only the initial error $\epsilon_{0}$; it does not compound.
C.4 Conclusion of Analysis
This model formally demonstrates that the traditional nested strategy leads to an exponential decay of the original context’s signal ($\alpha^{T}$), whereas FlowKV maintains the signal at a constant level of decay ($\alpha$). This provides a strong theoretical basis for why FlowKV is more effective at preserving long-range dependencies, supporting our empirical findings. The model also highlights that if the base compressor ($\mathcal{F}$) is exceedingly poor (i.e., $\alpha$ is very small), FlowKV’s benefits will be marginal. Conversely, with a hypothetical, near-lossless compression algorithm ($\alpha\approx 1$), the problem that FlowKV solves would become less significant.
Appendix D Dynamic Budget Allocation for Fair Comparison
A critical aspect of our experimental design is ensuring a fair comparison between FlowKV and baseline methods. This requires that at the end of any given turn, the total size of the KV cache managed by FlowKV is identical to the cache size of a baseline method using the same global compression ratio. We achieve this through a dynamic budget allocation mechanism, which we formalize below.
D.1 Step-by-Step Algorithm
At each conversational turn $t$, the per-turn compression rate is determined as follows:
1:Define Target Global Budget: Calculate the total target cache size for the current turn, $S_{target}(t)$, based on the size the full, uncompressed KV cache would have up to this turn, $S_{full}(t)$, and the fixed global compression ratio, $R_{global}$.
$S_{target}(t)=S_{full}(t)\times R_{global}$
2:Calculate Available Budget for New Data: Determine the budget available for the new KV pairs generated in the current turn ($Q_{t}$ and $R_{t}$), $B_{new}(t)$. This is done by subtracting the size of the already preserved cache from previous turns, $S_{preserved}(t-1)$, from the total target budget.
$B_{new}(t)=S_{target}(t)-S_{preserved}(t-1)$
3:Determine and Apply Local Compression: The newly generated KV pairs for turn $t$, which have an uncompressed size of $S_{new\_full}(t)$, are compressed to fit precisely into $B_{new}(t)$. This defines the local compression ratio for this specific turn, $R_{local}(t)$.
$R_{local}(t)=\frac{B_{new}(t)}{S_{new\_full}(t)}$
This local ratio is then applied to the new KV pairs using the underlying compression algorithm (e.g., SnapKV).
4:Update Preserved Cache: The resulting compressed KV pairs from this turn are appended to $S_{preserved}(t-1)$ to create the new, larger preserved cache, $S_{preserved}(t)$, that is carried to the next turn.
This dynamic, per-turn adjustment of the local compression rate ensures we strictly adhere to the global budget while isolating and preserving the integrity of previously compressed history.
D.2 Discussion on Aggressive Compression in Later Turns
A perceptive observation is that this dynamic approach may require more aggressive compression in later turns (i.e., $R_{local}(t)$ can be higher than $R_{global}$). This is a correct and integral aspect of FlowKV’s design.
This constitutes the core design trade-off of FlowKV: we trade potentially stronger, single-pass compression on new information for the perfect fidelity of historical information (by avoiding the cumulative information loss from re-compression). Our experimental results strongly demonstrate that this trade-off is extremely beneficial. For multi-turn conversations, avoiding catastrophic forgetting of early context (such as system prompts, user identity, and key facts) is far more important than minor variations in the local compression rate of later turns. By isolating historical information, FlowKV fundamentally solves the problem of information decay that plagues traditional methods as the number of turns accumulates.
Appendix E Evaluation Benchmark
E.1 Dataset Details